OpenStudy (anonymous):

Solve the following system of equations. -3x + 5y = 80 x + 5y = 40 A. x = 8, y = 6.4 B. x = -10, y = 10 C. x = 2, y = 17.2 D. x = 10, y = -10 @jojo4eva @iGreen @chosenmatt

OpenStudy (chosenmatt):

so? what do you think?

OpenStudy (anonymous):

A

OpenStudy (chosenmatt):

are u sure?

OpenStudy (igreen):

A is incorrect.. You can solve using substitution..first rearrange the 2nd one: x + 5y = 40 Subtract 5y to both sides: x = -5y + 40 Now plug in '-5y + 40' for 'x' in the 1st equation: -3x + 5y = 80 -3(-5y + 40) + 5y = 80 Distribute -3 into the parenthesis: 15y - 120 + 5y = 80 Combine like terms: 20y - 120 = 80 Add 120 to both sides, what's 120 + 80? @leeaah

OpenStudy (anonymous):

200 @iGreen

OpenStudy (igreen):

Yes, that gives us: 20y = 200 Now divide 20 to both sides, what's 200 / 20?

OpenStudy (anonymous):

-3x + 5y = 80 x + 5y = 40 You can do this a few ways. but one way would be to turn the +5y in the equation the opposite of each other so they cancel out and then solve for x. To do that what number can we times the equation by to get that? For -3x + 5y = 80 we can times it by -1 -1(-3x + 5y) = 80 *-1 3x + -5y = -80 Now we have 3x - 5y = -80 x + 5y = 40 Cancel out the 5, which will give us 3x = -80 x = 40 Can you finish to solve for x or do you still need help?

OpenStudy (igreen):

Can you divide 200 / 20?

OpenStudy (anonymous):

10

OpenStudy (igreen):

Yes, that gives us y = 10. And only one option has that..and what's that?

OpenStudy (anonymous):

d

OpenStudy (chosenmatt):

um no

OpenStudy (anonymous):

so its b

OpenStudy (chosenmatt):

yeah good job :)

OpenStudy (igreen):

Yes, B is correct.

OpenStudy (igreen):

D has y = -10, not 10. So it's incorrect.